# Numerical simulations of density perturbation and gravitational wave production from cosmological first-order phase transition

### v1

@@ IV NUMERICAL RESULTS
When vacuum bubbles expand and collide, the large amount of energy carried by their walls is released violently and creates asymmetric disturbances, making it a source of GWs Falkowski and No (2013); Konstandin and Servant (2011). In Fig. 5, we show the energy density spectrum of GW for FOPT. For $log_{10}[kR^{*}]\textless 0.5$, the GW power spectrum exhibits a steep rise with a slope of $k^{3}$, which is a little bit different for the $\alpha=0.5$ case and the discrepancy may be caused by the numerical fault. In the range of $0.8\textless log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-2}$, Moreover, our analysis reveals that the overall shape and amplitude of the power spectrum show little sensitivity to the parameters $\alpha$ and $\beta/H$. Meanwhile, according to our calculations for the $\alpha$ is $0.5$ and $1$, the magnitude is larger when $\beta/H$ is smaller, which would disppear at $\alpha=$ $5$ and $10$.
If this overdensity exceeds the threshold $\delta_{c}$, the Hubble patch collapses into a PBH Musco et al. (2005); Harada et al. (2013). During the the mid-to-late phase transition period, the value of threshold is within the range of 0.40 to 0.67 Musco (2019); Escrivà et al. (2020), in this work, we select commonly used $\delta_{c}=0.45$ Jedamzik and Niemeyer (1999); Green et al. (2004); Musco et al. (2005); Musco and Miller (2013); Harada et al. (2015). It is worth mentioning that the condition where the overdensity $\delta$ exceeds the critical value $\delta_{c}$ generally occurs before the vacuum energy fully decays in the overly dense regions. Furthermore, the results indicate that PBH formation occurs for slow phase transitions with $\beta/H=6$, see Appendix B and Appendix. C for more details during the PT processes. Refs. Liu et al. (2022); Kawana et al. (2023); Cai et al. (2024) suggest that a smaller $\beta/H$ leads to a higher PBH abundance, which aligns with our simulation results. Regarding the effect of $\alpha$, these studies conclude that PBH abundance increases as $\alpha$ increases. However, our simulation results indicate that $\alpha$ has little impact on PBH abundance, making it difficult to draw a definitive conclusion. In the future, with improved simulation precision, a more accurate assessment of $\alpha$’s effect may be possible.
V CONCLUSION AND DISCUSSION
@@
In this study, we performed three-dimensional numerical simulations to investigate the density perturbation and GWs production during FOPTs. Our simulations show that regions undergoing delayed vacuum decay can indeed develop local overdensities, and that smaller values of $\beta/H$ facilitate the formation of such overdense regions. In our simulations, variations in $\alpha$ did not significantly affect the formation of high-density regions, suggesting that higher-precision simulations may be needed to clarify this effect. Additionally, in the power spectrum of density perturbations at the completion of the phase transition, the slope follows $k^{3}$ for $\log_{10}[kR^{*}]\textless 0.5$, while in the range $0.8<\log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-1.5}$. We found that for $\alpha\textless 1$, $\delta_{H}$ is positively correlated with $\alpha$, which is consistent with the results given in Ref. Liu et al. (2023).
We observed two peaks in the density perturbation. The first peak is caused by the delay of vacuum decay in different Hubble volumes, while the second peak originates from the moving forward behavior of the bubble walls. Moreover, for $\alpha>1$, the peak due to the delay of vacuum decay dominates, whereas for $\alpha<1$, the peak arising from the moving forward behavior of the bubble walls becomes dominant. With the fitted curvature perturbation formula, for latter case with the same PT temperature, the CMB spectral distortion considered would yield more strong limit on the PT strength $\alpha$ and weaker limit on the PT inverse duration $\beta/H$ than given by Ref. Liu et al. (2023). The situation is in contrast with the strong PT scenarios of $\alpha>1$. For more detailed on the BBN and $\Delta N_{eff}$ constraints on the PT parameters, we refer to Refs.Bai and Korwar (2022); Deng and Bian (2023). For the superhorizon fluctuations induced by delayed vacuum decay at low scale PTs with $T_{PT}\lesssim$GeV, see Ref. Elor et al. (2024).
@@
Furthermore, the GW spectra were shown to have the slope following $k^{3}$ and $k^{-2}$ at small and large wave numbers. We did not observe a significant dependence on $\alpha$ and $\beta/H$ for these slope. Additionally, when analyzing the GW power spectrum at different stages of the phase transition, we did not find any notable changes in its slope. In the simulations with $\alpha=0.5$ and $\alpha=1$, we found that as $\beta/H$ decreases, the magnitude is larger. This finding is consistent with the conclusions of Ref. Lewicki et al. (2024).
@@ Acknowledgements.
Appendix A Phase transition process
The evolution of the scalar field $\phi$ is depicted in Fig. 6. It can be observed that the mean value of the field gradually increases from zero, corresponding to the nucleation and expansion of vacuum bubbles in space. Initially, the growth rate of the field’s mean value accelerates before slowing down and eventually reaching a plateau at zero. Moreover, as $\alpha$ increases and $\beta/H$ becomes larger, the completion time of the phase transition, characterized by $t/R^{*}$, decreases. This is expected, as a stronger phase transition facilitates a more rapid completion of the process. The time of phase transition completion and the corresponding number of bubbles are presented in Tab. 1.
Figure 6: The mean field evolution under weak phase transition strength (left, solid line $\alpha$=0.5, dashed line $\alpha$=1) and strong phase transition strength (right, solid line $\alpha$=5, dashed line $\alpha$=10). Four colors represent different $\beta/H$.
Table 1: Bubble parameters for different values of $\beta/H$ and $\alpha$. Each cell contains two quantities: the bubble number $N_{b}$ of the generated bubbles during the FOPT processes, the time $t/R^{*}$ required to reach $\phi/\phi_{b}$=0.999.
$\alpha\backslash\beta/H$ 6 8 10 12
$0.5$ $(281,3.57)$ $(391,4.12)$ $(521,4.67)$ $(601,5.02)$
$1$ $(296,2.71)$ $(397,3.15)$ $(498,3.56)$ $(610,3.93)$
$5$ $(289,1.83)$ $(388,2.13)$ $(481,2.42)$ $(601,2.72)$
$10$ $(278,1.70)$ $(387,2.01)$ $(500,2.30)$ $(598,2.54)$
As an illustration in Fig. 7, we present bubble slices with $\beta/H=6$ and $\alpha=5$, which visually demonstrate the process of bubble expansion and phase transition. Clearly, one could have more energy transormation from bubble walls’s moving forward with the proceeding of the PT. For more details on the energy evolutions of different components, see Fig. 8.
Figure 7: The bubble slice diagram for $\beta/H$ = 6 and $\alpha$ = 5. The dashed lines are presented to devide the simulation box to 64 Hubble patches.
Figure 8: The energy density $\rho$ under weak phase transition strength (left, $\alpha=1$, solid line $\beta/H=6$, dashed line $\beta/H=10$) and strong phase transition strength(right, $\alpha=10$ solid line $\beta/H=6$, dashed line $\beta/H=10$).
Appendix B Evolution of the density perturbation
Using the Press-Schechter formalismPress and Schechter (1974); Ando et al. (2018); Yoo et al. (2018), we obtain an alternative expression for $\delta_{H}(t)$
$\delta_{H}(t)=\int\frac{dk}{k}W^{2}(k)\mathcal{P_{\delta}}(t),$ (15)
Here, $W(k)$ is the window function, which in this case is chosen to be Gaussian
$W(k)=e^{-k^{2}/\mathcal{H}^{2}}\;.$ (16)
The results for different phase transition parameters are shown in Fig. 9, which demonstrate a quite similar behavior as in Fig. 2.
Figure 9: Evolution diagram of $\delta_{H}$ during FOPT process, with $\delta_{H}(t)=\int\frac{dk}{k}W^{2}(k)\mathcal{P_{\delta}}(t)$ where the contribution from smaller scales ($kR^{*}>1.3$) has been neglected. The $\alpha$ from left to right are 0.5, 1, 5, 10, respectively.
Figure 10: $\delta$ of the last 16 the last 16 delayed-decayed regions. The $\beta/H$ from left to right are 6, 8, 10, and 12, the $\alpha$ from top to bottom are 0.5, 1, 5, and 10, respectively.
We numerically analyze the evolution of local energy overdensities $\delta$ induced by delayed vacuum decay during FOPTs in the early universe. Specifically, we examine how different phase transition parameters, $\alpha$ and $\beta/H$, affect $\delta$ and whether these overdense regions can collapse into PBHs. As shown in the Fig. 10, due to the phase transition in the background universe, the vacuum energy gradually converts into the radiation energy, resulting in a lower energy density in the delay region compared to the background universe, and $\delta$ gradually decreases. When the delay region begins to undergo a phase transition, $\delta$ gradually increases. After the energy density disturbance reaches its maximum value, the content of the wall gradually decreases, and the background radiation energy is also averaged throughout the entire space, so $\delta$ gradually approaches zero. In addition, it can be observed that $\alpha$ has little effect on energy density perturbations, while its dependence on $\beta/H$ is extremely strong. The smaller the value of $\beta/H$, the larger the energy density perturbation. When $\beta/H$ = 6, the perturbation reaches the critical threshold $\delta_{c}$ = 0.45, ultimately collapsing to form PBHs. Our simulation results show that for larger $\beta/H$ (faster phase transitions), the probability of $\delta$ exceeding the critical threshold $\delta_{c}=0.45$ is significantly reduced. This trend suggests that PBH formation is suppressed in stronger and faster phase transitions. This phenomenon can be understood through the following mechanisms. When $\alpha$ is large, a greater fraction of vacuum energy is released during the phase transition, leading to a more uniform redistribution of energy and rapid smoothing of density fluctuations. As a result, compared to smaller $\alpha$ cases, the likelihood of forming sufficiently large $\delta$ to trigger PBH formation is reduced. Furthermore, when $\beta/H$ is large, the phase transition proceeds rapidly, with vacuum bubbles nucleating and merging within a short timescale. This results in a more homogeneous phase transition across space, reducing the formation of localized energy overdensities.
Appendix C GW power spectra
We compute the GW (GW) power spectrum generated during FOPTs and investigate its dependence on different values of the equation of state parameter $\omega$. Our results indicate that when $\alpha$ is 5 and 10, $\beta/H$ is 6 and 8, the slope of the GW power spectrum increases as omega increases, which do not change a lot for other circumstances. The GW power spectrum remains consistent in shape, with no significant variations observed. This suggests that the GW spectrum produced by bubble wall collisions during FOPTs is weakly dependent on $\omega$. However, we note that our conclusions regarding the dependence of the GW spectrum on $\omega$ differ from those reported in Refs. Domènech et al. (2020); Cai et al. (2020), requiring further investigation to clarify this issue.

### v2

@@ III Numerical results
When vacuum bubbles expand and collide, the large amount of energy carried by their walls is released violently and creates asymmetric disturbances, making it a source of GWs [Falkowski:2012fb, Konstandin:2011ds]. With the proceeding of PTs, the magnitudes of the GW spectra grow, as show in the APP. C. In Fig. 6, we show the energy density spectra of GWs for FOPTs. For $\log_{10}[kR_{*}]\textless 0.5$, the GW power spectrum exhibits a steep rise with a slope of $k^{3}$, which is a little bit different for the $\alpha=0.5$ case and the discrepancy may be caused by the numerical fault. In the range of $0.8\textless\log_{10}[kR_{*}]\textless 1.3$, the slope is $k^{-2}$, Moreover, our analysis reveals that the overall shape and amplitude of the power spectrum show little sensitivity to the parameters $\alpha$ and $\beta/H$. Meanwhile, according to our calculations for the $\alpha$ is $0.5$ and $1$, the magnitude is larger when $\beta/H$ is smaller, though this trend weakens for $\alpha=5$ and $10$.
We perform a fit to the gravitational‑wave spectrum in the presence of cosmic expansion, the fitting function is given below[Zhao:2022cnn],
$\Omega_{\text{gw}}=\tilde{\Omega}(k_{p})\frac{(a+b)^{c}k_{p}^{b}k^{a}}{\left(bk_{p}^{(a+b)/c}+ak^{(a+b)/c}\right)^{c}},$ (19)
we obtain $a=0.93,b=2.03,c=0.03$ from the fit. We adopt the parametrization of the GW spectrum given in Ref. [Cutting:2018tjt, Zhao:2022cnn]
$h^{2}\Omega_{\mathrm{GW}}(f)=h^{2}\tilde{\Omega}(f_{p})\frac{(a+b)^{c}f_{p}^{b}f^{a}}{\left(bf_{p}^{(a+b)/c}+af^{(a+b)/c}\right)^{c}},$ (20)
the magnitude of the low‑frequency GW spectrum is $\tilde{\Omega}(f_{p})=1.67\times 10^{-5}\left(\frac{100}{g_{\ast}(T_{\ast})}\right)^{1/3}\left(\frac{\alpha}{1+\alpha}\right)^{2}\left(\frac{H_{\ast}}{\beta}\right)^{2}$, with its low‑frequency peak located at $f_{p}=1.6\times 10^{-5}\frac{\beta}{H_{\ast}}\frac{T_{\ast}}{100}\left(\frac{g_{\ast}}{100}\right)^{1/6}\ \mathrm{Hz}$. In Fig. 7, we present the FOPT GW spectra obtained by substituting the fitted parameters into Eq. (20). The results also show cases detectable by LIGO [LIGOScientific:2014pky], Taiji [Ruan:2018tsw], and SKA [Carilli:2004nx].
Figure 7: GW spectra for three scenarios: ($\alpha=1,\beta/H=10,T=10^{-2}$ GeV), ($\alpha=1,\beta/H=10,T=10^{3}$ GeV), and ($\alpha=10,\beta/H=20,T=10^{7}$ GeV). The sensitivity curves including EPTA [EPTA:2015qep], PPTA [Shannon:2015ect], NANOGrav [NANOGRAV:2018hou], $\mu Ares$ [Sesana:2019vho], SKA [Carilli:2004nx], LISA [LISA:2017pwj], Taiji [Ruan:2018tsw], DECIGO [Kawamura:2011zz], LIGO [LIGOScientific:2014pky], CE [Reitze:2019iox], and ET [Punturo:2010zz].
IV Conclusion and discussion
@@
In this study, we performed three-dimensional numerical simulations to investigate the PBH abundance, density perturbation, and GWs production during FOPTs. Our simulations show that regions undergoing delayed vacuum decay can indeed develop local overdensities, and that smaller values of $\beta/H$ facilitate the formation of such overdense regions. Additionally, we computed the abundance of PBHs and found that it increases with $\alpha$ but decreases with $\beta/H$, which is consistent with the conclusions reported in Refs. [Liu:2021svg, Kawana:2022olo, Cai:2024nln]. Notably, the result exhibits stronger sensitivity to variations in $\beta/H$ and a relatively weaker dependence on $\alpha$. In particular, when $\alpha\gtrsim 10$, the influence of $\alpha$ on the PBH abundance becomes further suppressed.
In the power spectrum of density perturbations at the end of the phase transition, the slope follows $k^{3}$ for $\log_{10}[kR_{*}]\textless 0.5$, while in the range $0.8<\log_{10}[kR_{*}]\textless 1.3$, the slope is $k^{-1.5}$. We observed two peaks in the density perturbation. The first peak is caused by the delay of vacuum decay in different Hubble volumes, while the second peak originates from the moving forward behavior of the bubble walls. Moreover, for $\alpha<1$, the peak due to the delay of vacuum decay dominates, whereas for $\alpha>1$, the peak arising from the moving forward behavior of the bubble walls becomes dominant.
Since FOPT can generate significant curvature perturbations on small scales, the upper bound constraints on the power spectrum of curvature perturbations $\mathcal{P}_{\mathcal{R}}$ can be translated into limitations for the phase transition parameters $\alpha$ and $\beta/H$. Using the fitted curvature perturbation Eq. (18), our results yield less stringent constraints on $\alpha$ and $\beta/H$ compared to those reported in Ref. [Liu:2022lvz], but this result is slightly more stronger than that reported in Ref. [Cai:2024nln]. Such a PT is found to be in tension with the Hubble tension explanation [Elor:2023xbz]. For more detailed on the BBN and $\Delta N_{eff}$ constraints on the PT parameters, we refer to Refs. [Bai:2021ibt, Deng:2023twb].
@@
Furthermore, the GW spectra were shown to have the slope following $k^{3}$ and $k^{-2}$ at small and large wave numbers. We did not observe a significant dependence on $\alpha$ and $\beta/H$ for these slopes. Additionally, when analyzing the GW power spectrum at different stages of the phase transition, we did not find any notable changes in its slope. In the simulations with $\alpha=0.5$ and $\alpha=1$, we found that as $\beta/H$ decreases, the magnitude is larger. This finding is consistent with the conclusions of Ref. [Lewicki:2024ghw].
Appendix A Phase transition process
The evolution of the scalar field $\phi$ is depicted in Fig. 8. It can be observed that the mean value of the field gradually increases from zero, corresponding to the nucleation and expansion of vacuum bubbles in space. Initially, the growth rate of the field’s mean value accelerates before slowing down and eventually reaching a plateau. Moreover, as $\alpha$ increases and $\beta/H$ becomes larger, the completion time of the phase transition, characterized by $t/R_{*}$, decreases. This is expected, as a stronger phase transition facilitates a more rapid completion of the process. The time of phase transition completion and the corresponding number of bubbles are presented in Tab. 1.
Figure 8: The mean field evolution under weak phase transition strength (left, solid line $\alpha$=0.5, dash-dotted line $\alpha$=1) and strong phase transition strength (right, solid line $\alpha$=5, dash-dotted line $\alpha$=10). Four colors represent different $\beta/H$.
Table 1: Bubble parameters for different values of $\beta/H$ and $\alpha$. Each cell contains two quantities: the bubble number $N_{b}$ of the generated bubbles during the FOPT processes, the time $t/R_{*}$ required to reach $\phi/\phi_{b}$=0.999.
$\alpha\backslash\beta/H$ 6 8 10 12
$0.5$ $(281,3.57)$ $(391,4.12)$ $(521,4.67)$ $(601,5.02)$
11 $(296,2.71)$ $(397,3.15)$ $(498,3.56)$ $(610,3.93)$
55 $(289,1.83)$ $(388,2.13)$ $(481,2.42)$ $(601,2.72)$
1010 $(278,1.70)$ $(387,2.01)$ $(500,2.30)$ $(598,2.54)$
As an illustration in Fig. 9, we present bubble slices with $\beta/H=6$ and $\alpha=5$, which visually demonstrate the process of bubble expansion and phase transition. Clearly, one could have more energy transormation from bubble walls’s moving forward with the proceeding of the PT. Fig. 10 show the energy evolution of different components. The bubble wall energy initially increases, reaching a peak around $t/R_{\star}\approx 1$, after which it starts to decrease. As bubbles collide and merge, the energy of the bubble walls gradually converts into the background radiation energy.
Figure 9: The bubble slice diagram for $\beta/H$ = 6 and $\alpha$ = 5. The dashed lines are presented to devide the simulation box to 64 Hubble volumes.
Figure 10: The energy density $\rho$ under weak phase transition strength (left, $\alpha=1$, solid line $\beta/H=6$, dash-dotted line $\beta/H=10$) and strong phase transition strength (right, $\alpha=10$ solid line $\beta/H=6$, dash-dotted line $\beta/H=10$).
Appendix B Evolution of the density perturbation
Figure 11: $\delta$ of the last 16 delayed-decayed regions. The $\beta/H$ from left to right are 6, 8, 10, and 12, the $\alpha$ from top to bottom are 0.5, 1, 5, and 10, respectively.
We numerically analyze the evolution of local energy overdensities $\delta$ induced by delayed vacuum decay during FOPTs in the early universe. Specifically, we examine how different phase transition parameters, $\alpha$ and $\beta/H$, affect $\delta$ and whether these overdense regions can collapse into PBHs. As shown in the Fig. 11, due to the phase transition in the background universe, the vacuum energy gradually converts into the radiation energy, resulting in a lower energy density in the delay region compared to the background universe, and $\delta$ gradually decreases. When the delay region begins to undergo a phase transition, $\delta$ gradually increases. After the energy density disturbance reaches its maximum value, the content of the wall gradually decreases, and the background radiation energy is also averaged throughout the entire space, so $\delta$ gradually approaches zero. In addition, it can be observed that $\alpha$ has little effect on energy density perturbations, while its dependence on $\beta/H$ is extremely strong. The smaller the value of $\beta/H$, the larger the energy density perturbation. When $\beta/H$ = 6, the perturbation reaches the critical threshold $\delta_{c}$ = 0.45, ultimately collapsing to form PBHs.
Our simulation results show that for larger $\beta/H$ (faster phase transitions), the probability of $\delta$ exceeding the critical threshold $\delta_{c}=0.45$ is significantly reduced. This trend suggests that PBH formation is suppressed in faster phase transitions. This phenomenon can be understood through the following mechanisms. When $\beta/H$ is large, the phase transition proceeds rapidly, with vacuum bubbles nucleating and merging within a short timescale. This results in a more homogeneous phase transition across space, reducing the formation of localized energy overdensities.
Appendix C GW power spectra
We compute the GW (GW) power spectrum generated during FOPTs and investigate its dependence on different values of the equation of state parameter $\omega$. Our results indicate that when $\alpha$ is 5 and 10, $\beta/H$ is 6 and 8, the slope of the GW power spectrum increases as omega increases, which do not change a lot for other circumstances. The GW power spectrum remains consistent in shape, with no significant variations observed. This suggests that the GW spectrum produced by bubble wall collisions during FOPTs is weakly dependent on $\omega$. However, we note that our conclusions regarding the dependence of the GW spectrum on $\omega$ differ from those reported in Refs. [Domenech:2020kqm, Cai:2019cdl], requiring further investigation to clarify this issue.